# DiffAIL: Diffusion Adversarial Imitation Learning

### 11 Dec 2023 v1

@@ Abstract
Imitation learning aims to solve the problem of defining reward functions in real-world decision-making tasks. The current popular approach is the Adversarial Imitation Learning (AIL) framework, which matches expert state-action occupancy measures to obtain a surrogate reward for forward reinforcement learning. However, the traditional discriminator is a simple binary classifier and doesn’t learn an accurate distribution, which may result in failing to identify expert-level state-action pairs induced by the policy interacting with the environment. To address this issue, we propose a method named diffusion adversarial imitation learning (DiffAIL), which introduces the diffusion model into the AIL framework. Specifically, DiffAIL models the state-action pairs as unconditional diffusion models and uses diffusion loss as part of the discriminator’s learning objective, which enables the discriminator to capture better expert demonstrations and improve generalization. Experimentally, the results show that our method achieves state-of-the-art performance and significantly surpasses expert demonstration on two benchmark tasks, including the standard state-action setting and state-only settings. Our code can be available at an anonymous link https://github.com/ML-Group-SDU/DiffAIL.
@@ Introduction
To effectively tackle the aforementioned extrapolation error, Adversarial imitation learning (AIL) (Ho and Ermon 2016; Fu, Luo, and Levine 2017; Kostrikov et al. 2018; Ghasemipour, Zemel, and Gu 2020; Zhang et al. 2020, 2022) has become the most popular approach in imitation learning. Rather than minimizing the divergence between the expert policy and the agent policy, AIL employs online interactive learning to focus on minimizing the divergence between the joint state-action distribution induced by the expert policy and the learned policy. However, typical AIL methods often use a simplistic discriminator that does not learn a distribution, which may not accurately classify specific expert-level actions generated by the policy during interactions with the environment. This can result in the naive discriminator failing to distinguish expert-level behaviors generated by the policy, potentially hindering agent learning. Therefore, a powerful discriminator that can accurately capture the distribution of expert data is crucial for AIL.
@@
Although many studies (Song, Meng, and Ermon 2020; Lu et al. 2022) have focused on designing accelerated sampling algorithms for diffusion models to reduce the time cost of generating samples while maintaining high sample quality, we found that only a few tens of diffusion steps were sufficient to achieve excellent performance when modeling various environment tasks in MuJoCo, unlike image generation tasks that require thousands of sampling steps. Therefore, in our work, the primitive form of the diffusion model (Ho, Jain, and Abbeel 2020) is adequate and efficient.
@@
AIL can be viewed as a unified perspective to minimize f-divergences, including Gail (Ho, Jain, and Abbeel 2020), AIRL (Fu, Luo, and Levine 2017), FAIRL (Ghasemipour, Zemel, and Gu 2020), f-Gail (Zhang et al. 2020), etc. These methods aim to minimize divergence by matching the occupancy measure of expert demonstrations and policy data. The policy improvement process and the discriminator optimization process are similar to the GAN. However, these methods typically rely on on-policy sampling and often require millions or even tens of millions of interactions. To address this issue, DAC  (Kostrikov et al. 2018) extends Gail to the off-policy algorithm, significantly improving the learning efficiency of AIL methods and reducing the necessary online interaction to hundreds of thousands of steps. AEAIL (Zhang et al. 2022) proposes to use the reconstruction loss of Autoencoder as the reward function for AIL.
@@
  • Compared to naïve Gail, can DiffAIL provide a denser surrogate reward?
@@ D    Limitaion and future work
In this work, we introduced diffusion models into the adversarial imitation learning framework and proposed DiffAIL. Experiments show that this approach could significantly enhance adversarial imitation learning performance, owing to the powerful distribution-capturing capability of the diffusion discriminator. Although sampling inference is unnecessary during this adversarial training, computing the noise mse loss $Diff_{\phi}$ still requires multi-step noise sampling, which undoubtedly increases the computational cost of DiffAIL. However, our ablation studies on diffusion steps showed that more steps lead to better discriminator training, posing a trade-off between the two. Meanwhile, we also noticed that much research has focused on accelerating sampling(Song, Meng, and Ermon 2020; Bao et al. 2022; Lu et al. 2022; Liu, Gong, and Liu 2022; Liu et al. 2022; Song et al. 2023) in diffusion models - these methods can maintain generation quality while drastically reducing sampling steps, even enabling one-step generation(Liu, Gong, and Liu 2022; Song et al. 2023). Therefore, future Diffusion AIL could leverage advanced samplers to accelerate training and reduce computational costs.

### 12 Dec 2023 v2

@@ Abstract
Imitation learning aims to solve the problem of defining reward functions in real-world decision-making tasks. The current popular approach is the Adversarial Imitation Learning (AIL) framework, which matches expert state-action occupancy measures to obtain a surrogate reward for forward reinforcement learning. However, the traditional discriminator is a simple binary classifier and doesn’t learn an accurate distribution, which may result in failing to identify expert-level state-action pairs induced by the policy interacting with the environment. To address this issue, we propose a method named diffusion adversarial imitation learning (DiffAIL), which introduces the diffusion model into the AIL framework. Specifically, DiffAIL models the state-action pairs as unconditional diffusion models and uses diffusion loss as part of the discriminator’s learning objective, which enables the discriminator to capture better expert demonstrations and improve generalization. Experimentally, the results show that our method achieves state-of-the-art performance and significantly surpasses expert demonstration on two benchmark tasks, including the standard state-action setting and state-only settings. Our code can be available at the link https://github.com/ML-Group-SDU/DiffAIL.
@@ Introduction
To effectively tackle the aforementioned extrapolation error, Adversarial imitation learning (AIL) (Ho and Ermon 2016; Fu, Luo, and Levine 2017; Kostrikov et al. 2018; Ghasemipour, Zemel, and Gu 2020; Zhang et al. 2020; Garg et al. 2021; Zhang et al. 2022) has become the most popular approach in imitation learning. Rather than minimizing the divergence between the expert policy and the agent policy, AIL employs online interactive learning to focus on minimizing the divergence between the joint state-action distribution induced by the expert policy and the learned policy. However, typical AIL methods often use a simplistic discriminator that does not learn a distribution, which may not accurately classify specific expert-level actions generated by the policy during interactions with the environment. This can result in the naive discriminator failing to distinguish expert-level behaviors generated by the policy, potentially hindering agent learning. Therefore, a powerful discriminator that can accurately capture the distribution of expert data is crucial for AIL.
@@
Although many studies (Song, Meng, and Ermon 2020; Lu et al. 2022a, b; Zheng et al. 2023) have focused on designing accelerated sampling algorithms for diffusion models to reduce the time cost of generating samples while maintaining high sample quality, we found that only a few tens of diffusion steps were sufficient to achieve excellent performance when modeling various environment tasks in MuJoCo, unlike image generation tasks that require thousands of sampling steps. Therefore, in our work, the primitive form of the diffusion model (Ho, Jain, and Abbeel 2020) is adequate and efficient.
@@
AIL can be viewed as a unified perspective to minimize f-divergences, including Gail (Ho, Jain, and Abbeel 2020), AIRL (Fu, Luo, and Levine 2017), FAIRL (Ghasemipour, Zemel, and Gu 2020), f-Gail (Zhang et al. 2020), etc. These methods aim to minimize divergence by matching the occupancy measure of expert demonstrations and policy data. The policy improvement process and the discriminator optimization process are similar to the GAN (Goodfellow et al. 2014). However, these methods typically rely on on-policy sampling and often require millions or even tens of millions of interactions. To address this issue, DAC  (Kostrikov et al. 2018) extends Gail to the off-policy algorithm, significantly improving the learning efficiency of AIL methods and reducing the necessary online interaction to hundreds of thousands of steps. AEAIL (Zhang et al. 2022) proposes to use the reconstruction loss of Autoencoder as the reward function for AIL.
@@
  • Compared to naïve Gail, can DiffAIL provide a more linearly correlational surrogate reward?
@@ Acknowledgments
This work was supported by the NSFC (Nos. 62206159, 62176139), the Natural Science Foundation of Shandong Province (Nos. ZR2022QF117, ZR2021ZD15), the Fundamental Research Funds of Shandong University, the Fundamental Research Funds for the Central Universities. G. Wu was sponsored by the TaiShan Scholars Program.
@@ D    Limitaion and future work
In this work, we introduced diffusion models into the adversarial imitation learning framework and proposed DiffAIL. Experiments show that this approach could significantly enhance adversarial imitation learning performance, owing to the powerful distribution-capturing capability of the diffusion discriminator. Although sampling inference is unnecessary during this adversarial training, computing the noise mse loss $Diff_{\phi}$ still requires multi-step noise sampling, which undoubtedly increases the computational cost of DiffAIL. However, our ablation studies on diffusion steps showed that more steps lead to better discriminator training, posing a trade-off between the two. Meanwhile, we also noticed that much research has focused on accelerating sampling(Song, Meng, and Ermon 2020; Bao et al. 2022; Lu et al. 2022a; Liu, Gong, and Liu 2022; Liu et al. 2022; Song et al. 2023) in diffusion models - these methods can maintain generation quality while drastically reducing sampling steps, even enabling one-step generation(Liu, Gong, and Liu 2022; Song et al. 2023). Therefore, future Diffusion AIL could leverage advanced samplers to accelerate training and reduce computational costs.